OpenStudy (anonymous):

8x – 3y = 1 –2x + 3y = 11 A.(–1, –3) B.(–1, 3) C.(2, 5) D.(5, 2)

OpenStudy (luigi0210):

Add up both equations

OpenStudy (shamil98):

8x - 3y = 1 -2x +3y = 11 y crosses up so -6x = 12 divide both sides by -6 and you have your answer.

OpenStudy (anonymous):

tanka you c:

OpenStudy (anonymous):

wait the 2 is for the x or y?

OpenStudy (shamil98):

6x = 12 x =2 i accidentally put -6x my bad. the only option with 2 as x is C.

OpenStudy (anonymous):

omg thankks. c: